OpenStudy (omarbirjas):

z varies jointly with x and y. when x=2 and y=2 z=7. what is z when x=4 and y=8. Please show work...

OpenStudy (omarbirjas):

@myininaya @nincompoop

OpenStudy (omarbirjas):

@dan815

OpenStudy (omarbirjas):

@TheSmartOne @Luigi0210

OpenStudy (omarbirjas):

@BloomLocke367

OpenStudy (omarbirjas):

@confluxepic PLZZ

OpenStudy (confluxepic):

Let me think about it. @omarbirjas

OpenStudy (luigi0210):

*waits on Conflux*

OpenStudy (confluxepic):

The equation would be z=kxy.

OpenStudy (confluxepic):

K is the constant.

OpenStudy (confluxepic):

To turn the problem into an equation you have to add k.

OpenStudy (confluxepic):

Plug in the values.

OpenStudy (omarbirjas):

so what would the constant be? any number i wish or what?

OpenStudy (omarbirjas):

z=k4*8

OpenStudy (confluxepic):

You first solve for the constant.

OpenStudy (omarbirjas):

ohhh so we use the previous values.

OpenStudy (omarbirjas):

k=2*2*7

OpenStudy (confluxepic):

7=k(2)(2)

OpenStudy (confluxepic):

Find k and tell me.

OpenStudy (omarbirjas):

k=28

OpenStudy (confluxepic):

Can you tell me how you got that?

OpenStudy (omarbirjas):

7*2=14

OpenStudy (omarbirjas):

14*2=28

OpenStudy (confluxepic):

\(\color{blue}{\text{Originally Posted by}}\) @confluxepic 7=k(2)(2) \(\color{blue}{\text{End of Quote}}\) This is the equation.

OpenStudy (confluxepic):

It would be k=7/4.

OpenStudy (omarbirjas):

ohhhh yea i didnt do the deviding part, i just swapped out k and 7

OpenStudy (omarbirjas):

ok now we can use z=(7/4)*(4)*(8)to find z right?

OpenStudy (confluxepic):

Now you have to find z when x=4 and y=8.

OpenStudy (confluxepic):

Yes.

OpenStudy (omarbirjas):

z=56

OpenStudy (omarbirjas):

god bless you and wolframalpha.com

OpenStudy (omarbirjas):

btw cute cat......and baby at @Nnesha

OpenStudy (confluxepic):

Thanks. :)

OpenStudy (confluxepic):

Yes. z=56.